OpenStudy (anonymous):

Let A be the area of a circle with radius r. If dr/dt=4, find dA/dt when r=3

OpenStudy (anonymous):

Area of a circle is A = pi * r^2 You have to find dA/dt, so just differentiate the function above with respect to t

OpenStudy (anonymous):

The formula for the area of a circle is A = pi * r^2 or A = 3.14 * r squared.

OpenStudy (anonymous):

2(3.14)r?

OpenStudy (anonymous):

I am sorry, but I am completely lost in this subject so I don't know what I am doing at all

OpenStudy (anonymous):

Differentiate the area function with respect to t: d/dt A = d/dt [pi * r^2] I hope you are familiar with the chain rule, because we apply it in the next step to get: dA/dt = 2 pi r dr/dt Now just plug in the values of r and dr/dt, dA/dt = 2 * pi * 3 * 4 = 24 pi

OpenStudy (anonymous):

great

OpenStudy (anonymous):

Ok so pi (pronounced pie) is basically just 3.14. So if I say 5*(times) pi, I mean 5 * 3.14. See? The radius of a circle is just half of the diameter of the circle. If I say that the radius is squared then I mean the radius times the radius. If the radius is 5 then it would be 5 times itself so 5 * 5.

OpenStudy (anonymous):

@fbabygrl let me know if you did not get any step in my answer posted above.